# Fully Adaptive Zeroth-Order Method for Minimizing Functions with Compressible Gradients

### v2

@@ Abstract
We propose an adaptive zeroth-order method for minimizing differentiable functions with $L$-Lipschitz continuous gradients. The method is designed to take advantage of the eventual compressibility of the gradient of the objective function, but it does not require knowledge of the approximate sparsity level $s$ or the Lipschitz constant $L$ of the gradient. We show that the new method performs no more than $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$ function evaluations to find an $\epsilon$-approximate stationary point of an objective function with $n$ variables. Assuming additionally that the gradients of the objective function are compressible, we obtain an improved complexity bound of $\mathcal{O}\left(s\log\left(n\right)\epsilon^{-2}\right)$ function evaluations, which holds with high probability. Preliminary numerical results illustrate the efficiency of the proposed method and demonstrate that it can significantly outperform its non-adaptive counterpart. Accompanying code can be found at https://github.com/mines-opt-ml/ZORO-FA.
Keywords: Derivative-Free Optimization; Black-Box Optimization; Zeroth-Order Optimization; Worst-Case Complexity; Compressible Gradients
@@ 1 Introduction
In this work we consider the unconstrained minimization of a possibly nonconvex differentiable function $f:\mathbb{R}^{n}\to\mathbb{R}$. Most well-known iterative algorithms for this type of problem are gradient-based, meaning at each iteration they require at least one evaluation of the gradient of $f(\,\cdot\,)$. However, in many problems, it is impossible to compute the gradient of the underlying objective function. Examples include simulation-based optimization [18], reinforcement learning [22, 15, 8] and hyperparameter tuning [3, 25]. These settings create a demand for zeroth-order optimization algorithms, which are algorithms that rely only on evaluations of the objective function $f(\,\cdot\,)$. Recently, problems requiring zeroth-order methods have emerged in which $n$, the dimension of the problem, is of order $10^{5}$ or greater. In particular, we highlight many recent works applying zeroth-order methods to fine-tune Large Language Models (LLMs) [14, 28, 13, 7]. For problems of this scale, classical zeroth-order algorithms can struggle to make progress as they generically require $\mathcal{O}(n)$ function evaluations (“queries”) per iteration. As such, various structural assumptions on $\nabla f(x)$ have recently been proposed which allow for robust algorithms making as few as $\mathcal{O}(\log n)$ queries per iteration with high probability.
@@
Here, we focus on methods that exploit (approximate) gradient sparsity [26, 1, 4, 5, 6, 29], particularly the ZORO-type algorithms [4, 6] which use tools from signal processing to robustly approximate $\nabla f(x)$. One adverse consequence of this fusion of optimization and signal processing techniques is that ZORO-type algorithms have an unusually large number of hyperparameters. In this paper, we address this issue by proposing a fully adaptive version of ZORO, which we call ZORO-FA. The proposed method is designed to take advantage of the eventual compressibility of the gradient of the objective function, in which case a suitable gradient approximation can be obtained with high probability using $\mathcal{O}\left(s\log(n)\right)$ function evaluations. Specifically, at each iteration, the new method first attempts to use these cheap gradient approximations, and only when they fail is a forward-finite difference gradient computed using $\mathcal{O}\left(n\right)$ function evaluations. Moreover, ZORO-FA is endowed with a novel line-search procedure for selecting multiple hyperparameters at once, adapted from that proposed in [11]. We show that the new method takes at most $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$ function evaluations to find $\epsilon$-approximate stationary points. Assuming that the gradients of the objective function are compressible with sparsity level $s$, we obtain an improved complexity bound of $\mathcal{O}\left(s\log\left(n\right)\epsilon^{-2}\right)$ function evaluations, which holds with high probability. Our preliminary numerical results indicate that ZORO-FA requires significantly fewer function evaluations than ZORO to find points with similar function values, and that the improvement factor increases as the dimensionality of the problems grows.
@@ 2 Auxiliary Results
    where $|u|_{(j)}$ denotes the $i$-th largest-in-magnitude component of $u\in\mathbb{R}^{n}$.
@@
$\left[\nabla f(x)\right]_{(s)}=\arg\min\left\{\|v-\nabla f(x)\|_{2}\,:\,\|v\|_{0}\leq s\right\}.$
Notice that $\left[\nabla f(x)\right]_{(s)}$ is the $s$-sparse vector formed by the $s$ largest (in modulus) entries of $\nabla f(x)$.
Lemma 2.1.
@@
Proof.
See Section 2.5 in [19].. ∎
@@
Lemma 2.2.
(Proposition 3.5 in [19]) If $Z\in\mathbb{R}^{m\times n}$ satisfies the $4s$-RIP for some $s\in\left\{1,\ldots,n\right\}$, then for any $v\in\mathbb{R}^{n}$ we have
@@
Lemma 2.3.
(Theorem 5.2 in [2]) Given $s\in\mathbb{N}$ such that $m:=\lceil bs\ln(n)\rceil<n$, let $Z\in\mathbb{R}^{m\times n}$ be defined as
@@
$b>\dfrac{4\left(1+\dfrac{1+\ln(\frac{12}{\delta})}{\ln(\frac{n}{4s})}\right)}{c_{0}(\frac{\delta}{2})},$ (2)
for $c_{0}(\,\cdot\,)$ defined in (1), then $Z$ satisfies the $4s$-RIP for $\delta_{4s}(Z)=\delta$, with probability $\geq 1-2e^{-\gamma m}$, where
$\gamma=\left[c_{0}\left(\frac{\delta}{2}\right)-\frac{4}{b}\left(1+\dfrac{1+\ln(\frac{12}{\delta})}{\ln(\frac{n}{4s})}\right)\right].$
@@ 2.1 Gradient Approximation
ZORO [6] uses finite differences along random directions to approximate directional derivatives, thought of as linear measurements of $\nabla f(x)$. More precisely, let $\left\{z_{i}\right\}_{i=1}^{m}\subset\mathbb{R}^{n}$ be Rademacher random vectors. For fixed $h>0$ define the vector $y\in\mathbb{R}^{m}$ with
@@
where $\|w\|\leq\frac{Ln}{2}$.
Proof.
@@
We view (5) as a linear inverse problem to be solved for $\nabla f(x)$. As in [6], given $s\in\mathbb{N}\setminus\left\{0\right\}$, we will compute an approximation to $\nabla f(x)$ by approximately solving the problem
$\min\|Zv-y\|\quad\text{s.t.}\quad\|v\|_{0}\leq s,$ (6)
using the algorithm CoSaMP proposed in [19].
Lemma 2.5.
@@
with $\|w\|_{2}\leq Ln/2$. Thus, by Theorem 5 and Remark 3 in [9] we have
@@
Then, using Lemma 2.1, we obtain
@@
From Lemmas 2.1 and 2.2, and the assumption $\delta_{4s}(Z)\leq 0.22665$, we also have
@@
which implies that (7) is true.
The following corollary may be deduced from Lemma 2.5 by direct computation.
Corollary 2.6.
Let $\theta,\epsilon\in(0,1)$. Under the assumptions of Lemma 2.5, if
$\ell=\left\lceil\dfrac{\log(\theta/4)}{\log(0.5)}\right\rceil,$ (10)
@@
We call the quantity $s(\theta,p)$ defined in (11) the effective sparsity level.
@@ 2.2 Sufficient Decrease Condition
From Corollary 2.6, we see that if $s$ is sufficiently large and $h$ is sufficiently small, then CoSaMP applied to (6) is able to find a vector $g$ that satisfies
@@
Lemma 2.8.
Suppose that A1 holds. Given $x\in\mathbb{R}^{n}$ and $\epsilon>0$, let $g\in\mathbb{R}^{n}$ be given by
@@ 3 New Method and its Complexity
Step 0. Given $x_{0}\in\mathbb{R}^{n}$, $\epsilon\in(0,1)$, $\theta\in(0,1/2)$, $b\geq 1$, and $\sigma_{0}>0$. Choose $s_{0}$ such that $\lceil bs_{0}\ln(n)\rceil\leq n$, sample i.i.d. Rademacher random vectors $\{z_{i}\}_{i=1}^{n}$, and set $k:=0$.
@@
$\left[y^{(j)}_{k}\right]_{i}=\dfrac{f(x_{k}+h_{j}z_{i}^{(j)})-f(x_{k})}{h_{j}\sqrt{m_{j}}},\quad i=1,\ldots,m_{j}.$
@@
Remark 3.1.
Note that ZORO-FA bears some resemblence to the adaptive ZORO, or adaZORO algorithm presented in [6]. However, that algorithm only selects the target sparsity ($s$) adaptively. Moreover, we are able to provide theoretical complexity guarantees for ZORO-FA, whereas no such guarantees are presented for adaZORO.
@@ 3.1 Oracle Complexity Without Compressible Gradients
Lemma 3.2.
@@
In view of (27), (30) and assumption $\|\nabla f(x_{k})\|>\epsilon$, it follows from Lemma 2.8 that (14) holds for $x=x_{k}$ and $g=g_{k}^{(j_{k}-1)}$. Thus, by (29) and Lemma 2.7 we would have
@@
As a consequence of Lemma 3.2, we have the following result.
Lemma 3.3.
@@
Summing up these inequalities and using A2 and Lemma 3.3, we obtain
@@
Proof.
Let us consider the $k$th iteration of Algorithm 1. For any $j\in\left\{0,\ldots,j_{k}\right\}$, the computation of $g_{k}^{(j)}$ requires at most $n+1$ function evaluations. Moreover, checking the acceptance condition (24) at Step 2.5 requires one additional function evaluation. Thus, the total number of function evaluations at the $k$th iteration is bounded from above by $(n+1)(j_{k}+1)$. Therefore, by Lemma 3.2,
@@
Remark 3.6.
Recall that $m_{0}=bs_{0}\log(n)$. Thus, it follows from Corollary 3.5 that
@@
Therefore, when $m_{0}=n$, then Algorithm 1 requires no more than $\mathcal{O}\left(n\epsilon^{-2}\right)$ function evaluations to find an $\epsilon$-approximate stationary point of $f(\,\cdot\,)$. Regarding the dependence on $n$ and $\epsilon$, this bound is consistent,in order, with the complexity bounds established in [10, 11]. However, to take advantage of potential compressibility of gradients, one should take $m_{0}<<n$. In this case, the worst-case complexity bound becomes $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$.

### v3

@@ Abstract
We propose an adaptive zeroth-order method for minimizing differentiable functions with $L$-Lipschitz continuous gradients. The method is designed to take advantage of the eventual compressibility of the gradient of the objective function, but it does not require knowledge of the approximate sparsity level $s$ or the Lipschitz constant $L$ of the gradient. We show that the new method performs no more than $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$ function evaluations to find an $\epsilon$-approximate stationary point of an objective function with $n$ variables. Assuming additionally that the gradients of the objective function are compressible, we obtain an improved complexity bound of $\mathcal{O}\left(s\log\left(n\right)\epsilon^{-2}\right)$ function evaluations, which holds with high probability. Preliminary numerical results illustrate the efficiency of the proposed method and demonstrate that it can significantly outperform its non-adaptive counterpart.
Keywords:
derivative-free optimization black-box optimizationzeroth-order optimization worst-case complexity compressible gradients
@@ 1 Introduction
In this work we consider the unconstrained minimization of a possibly nonconvex differentiable function $f:\mathbb{R}^{n}\to\mathbb{R}$. Most well-known iterative algorithms for this type of problem are gradient-based, meaning at each iteration they require at least one evaluation of the gradient of $f(\,\cdot\,)$. However, in many problems, it is impossible to compute the gradient of the underlying objective function. Examples include simulation-based optimization Nakamura2017 , reinforcement learning salimans2017evolution ; mania2018simple ; choromanski2020provably and hyperparameter tuning bergstra2012random ; tett2022calibrating . These settings create a demand for zeroth-order optimization algorithms, which are algorithms that rely only on evaluations of the objective function $f(\,\cdot\,)$. Recently, problems requiring zeroth-order methods have emerged in which $n$, the dimension of the problem, is of order $10^{5}$ or greater. In particular, we highlight many recent works applying zeroth-order methods to fine-tune Large Language Models (LLMs) malladi2024fine ; zhang2024revisiting ; liu2024sparse ; chen2023deepzero . For problems of this scale, classical zeroth-order algorithms can struggle to make progress as they generically require $\mathcal{O}(n)$ function evaluations (“queries”) per iteration. For problems of moderate to large scale, deterministic zeroth-order algorithms often struggle to make progress, as they typically require $\mathcal{O}(n)$ function evaluations (or “queries”) per iteration. To mitigate this issue, many randomized zeroth-order methods have been proposed in recent years, with cheaper per-iteration cost in terms of function evaluations. Some of these methods mimic classical zeroth-order methods, but applied in a low-dimensional subspace Nesterov ; Bergou ; Kozak ; Cartis ; Roberts . Others aim to exploit ‘low-dimensional’ structure within $f$ Bandeira ; cartis2022dimensionality ; cartis2023bound ; wang2016bayesian ; Yue2023zeroth .
@@
Here, we focus on methods that exploit a certain kind of low dimensional structure, namely (approximate) gradient sparsity wang2018stochastic ; balasubramanian2018zeroth ; cai2021zeroth ; cai2022one ; cai2022zeroth ; qiu2024gradient , particularly the ZORO-type algorithms cai2021zeroth ; cai2022zeroth which robustly approximate $\nabla f(x)$ using tools from compressed sensing [^1] In Bandeira , compressed sensing techniques have been used to exploit (approximate) sparsity of $\nabla^{2}f(x)$ in the context of derivative-free trust-region methods.. One adverse consequence of this fusion of optimization and signal processing techniques is that ZORO-type algorithms have an unusually large number of hyperparameters. In this paper, we address this issue by proposing a fully adaptive version of ZORO, which we call ZORO-FA. The proposed method is designed to take advantage of the eventual compressibility of the gradient of the objective function, in which case a suitable gradient approximation can be obtained with high probability using $\mathcal{O}\left(s\log(n)\right)$ function evaluations. Specifically, at each iteration, the new method first attempts to use these cheap gradient approximations, and only when they fail a forward-finite difference gradient is computed using $\mathcal{O}\left(n\right)$ function evaluations. Moreover, ZORO-FA is endowed with a novel line-search procedure for selecting multiple hyperparameters at once, adapted from that proposed in grapiglia2024quadratic . We show that the new method takes at most $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$ function evaluations to find $\epsilon$-approximate stationary points. Assuming that the gradients of the objective function are compressible with sparsity level $s$, we obtain an improved complexity bound of $\mathcal{O}\left(s\log\left(n\right)\epsilon^{-2}\right)$ function evaluations, which holds with high probability. Our preliminary numerical results indicate that ZORO-FA requires significantly fewer function evaluations than ZORO to find points with similar function values, and that the improvement factor increases as the dimensionality of the problems grows.
@@ 2 Auxiliary Results
    where $|u|_{(j)}$ denotes the $j$-th largest-in-magnitude component of $u\in\mathbb{R}^{n}$.
@@
$\left[\nabla f(x)\right]_{(s)}=\arg\min\left\{\|v-\nabla f(x)\|_{2}\,:\,\|v\|_{0}\leq s\right\},$
where $\|v\|_{0}$ denotes the cardinality of the support of $v$. Note that $\left[\nabla f(x)\right]_{(s)}$ is the $s$-sparse vector formed by the $s$ largest (in modulus) entries of $\nabla f(x)$.
Lemma 1
@@
Proof
See Section 2.5 in needell2009cosamp .∎.
@@
Lemma 2
(Proposition 3.5 in needell2009cosamp ) If $Z\in\mathbb{R}^{m\times n}$ satisfies the $4s$-RIP for some $s\in\left\{1,\ldots,n\right\}$, then for any $v\in\mathbb{R}^{n}$ we have
@@
Lemma 3
(Theorem 5.2 in baraniuk2008simple ) Given $s\in\mathbb{N}$ such that $m:=\lceil bs\ln(n)\rceil<n$, let $Z\in\mathbb{R}^{m\times n}$ be defined as
@@
$b>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c_{1}(\delta)\equiv}\dfrac{4\left(1+\dfrac{1+\ln(\frac{12}{\delta})}{\ln(\frac{n}{4s})}\right)}{c_{0}(\frac{\delta}{2})},$ (2)
for $c_{0}(\,\cdot\,)$ defined in (1), then $Z$ satisfies the $4s$-RIP for $\delta_{4s}(Z)=\delta$, with probability $\geq 1-2e^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\gamma(\delta)}m}$, where
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\gamma(\delta)\equiv}\left[c_{0}\left(\frac{\delta}{2}\right)-\frac{4}{b}\left(1+\dfrac{1+\ln(\frac{12}{\delta})}{\ln(\frac{n}{4s})}\right)\right].$ (3)
@@ 2.1 Gradient Approximation
ZORO cai2022zeroth uses finite differences along random directions to approximate directional derivatives, thought of as linear measurements of $\nabla f(x)$. More precisely, let $\left\{z_{i}\right\}_{i=1}^{m}\subset\mathbb{R}^{n}$ be Rademacher random vectors. For fixed $h>0$ define the vector $y\in\mathbb{R}^{m}$ with
@@
where $w\in\mathbb{R}^{m}$ with $\|w\|\leq\frac{Ln}{2}$.
Proof
@@
We view (6) as a linear inverse problem to be solved for $\nabla f(x)$. As in cai2022zeroth , given $s\in\mathbb{N}\setminus\left\{0\right\}$, we will compute an approximation to $\nabla f(x)$ by approximately solving the problem
$\min_{v\in\mathbb{R}^{n}}\|Zv-y\|\quad\text{s.t.}\quad\|v\|_{0}\leq s,$ (7)
using the algorithm CoSaMP proposed in needell2009cosamp .
Lemma 5
@@
with $\|w\|_{2}\leq Ln/2$. Thus, by Theorem 5 and Remark 3 in foucart2012sparse we have
@@
Then, using Lemma 1, we obtain
@@
From Lemmas 1 and 2, and the assumption $\delta_{4s}(Z)<0.22665$, we also have
@@
which implies that (8) is true.∎
The following corollary may be deduced from Lemma 5 by direct computation.
Corollary 1
Let $\theta,\epsilon\in(0,1)$. Under the assumptions of Lemma 5, if
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ell\geq}\left\lceil\dfrac{\log(\theta/4)}{\log(0.5)}\right\rceil,$ (11)
@@
We call the quantity $s(\theta,p)$ defined in (12) the effective sparsity level. We emphasize that $s(\theta,p)$ is independent of $n$.
Note that, from Lemma 5, the number of iterations of CoSaMP only affects the factor $(1/2)^{\ell}$ in (8), leaving the other factors unchanged. The per-iteration cost of CoSaMP is $\mathcal{O}(mn+ms^{2})$ operations. Thus, Corollary 1 suggests that it is sufficient to perform
$\ell=\left\lceil\dfrac{\log(\theta/4)}{\log(0.5)}\right\rceil$
iterations to guarantee (14), provided that (12) and (13) hold, thereby saving computational effort. For example, if $\theta=0.25$, it follows that $\ell=4$ iterations of CoSaMP would be enough.
@@ 2.2 Sufficient Decrease Condition
From Corollary 1, we see that if $s$ is sufficiently large and $h$ is sufficiently small, then CoSaMP applied to (7) is able to find a vector $g$ that satisfies
@@
Lemma 7
Suppose that A1 holds. Given $x\in\mathbb{R}^{n}$, $\epsilon,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\theta}>0$, let $g\in\mathbb{R}^{n}$ be given by
@@ 3 New Method and its Complexity
Step 0. Given $x_{0}\in\mathbb{R}^{n}$, $\epsilon\in(0,1)$, $\theta\in(0,1/2)$, $b\geq 1$, and $\sigma_{0}>0$. Choose $s_{0}\in\mathbb{N}$ such that $\lceil bs_{0}\ln(n)\rceil{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq n/4}$, sample i.i.d. Rademacher random vectors $\{z_{i}\}_{i=1}^{n}$, and set $k:=0$.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left[y^{(j)}_{k}\right]_{i}=\dfrac{f(x_{k}+h_{j}z_{i})-f(x_{k})}{\sqrt{m_{j}}h_{j}},\quad i=1,\ldots,m_{j}}.$
@@
Remark 1
Note that vectors $\left\{z_{i}\right\}_{i=1}^{n}$ are sampled only once, at Step 0. Moreover, in step 2.2 it is important to evaluate the finite difference using the unscaled $z_{i}$ (i.e. compute $f(x_{k}+hz_{i})$), not the $i$-th column of $Z^{(j)}$.
Remark 2
Note that ZORO-FA bears some resemblence to the adaptive ZORO, or adaZORO algorithm presented in cai2022zeroth . However, that algorithm only selects the target sparsity ($s$) adaptively. Moreover, we are able to provide theoretical complexity guarantees for ZORO-FA, whereas no such guarantees are presented for adaZORO.
@@ 3.1 Oracle Complexity Without Compressible Gradients
In this subsection, we analyze Algorithm 1 under Assumptions A1 and A2 only; that is, we do not assume that the gradients of $f(\,\cdot\,)$ are compressible. We start by establishing an upper bound on the number of inner loops $j_{k}$ required to find a point that yields (25).
Lemma 8
@@
In view of (28), (31) and assumption $\|\nabla f(x_{k})\|>\epsilon$, it follows from Lemma 7 that (15) holds for $x=x_{k}$ and $g=g_{k}^{(j_{k}-1)}$. Thus, by (30) and Lemma 6 we would have
@@
As a consequence of Lemma 8, we have the following result.
Lemma 9
@@
Summing up these inequalities and using A2 and Lemma 9, we obtain
@@
Proof
Let us consider the $k$th iteration of Algorithm 1. For any $j\in\left\{0,\ldots,j_{k}\right\}$, the computation of $g_{k}^{(j)}$ requires at most $n+1$ function evaluations. Moreover, checking the acceptance condition (25) at Step 2.5 requires one additional function evaluation. Thus, the total number of function evaluations at the $k$th iteration is bounded from above by $(n+1)(j_{k}+1)$. Therefore, by Lemma 8,
@@
Remark 3
Recall that $m_{0}=\left\lceil bs_{0}\log(n)\right\rceil$. Thus, it follows from Corollary 2 that
@@
Therefore, our complexity bound ranges from $\mathcal{O}(n\epsilon^{-2})$ to $\mathcal{O}(n^{2}\epsilon^{-2})$, depending on the choice of $s_{0}$. When $s_{0}$ is chosen such that $m_{0}=\left\lceil bs_{0}\ln(n)\right\rceil=n/4$, Algorithm 1 requires no more than $\mathcal{O}\left(n\epsilon^{-2}\right)$ function evaluations to find an $\epsilon$-approximate stationary point of $f(\,\cdot\,)$. Regarding the dependence on $n$ and $\epsilon$, this bound is consistent, in order, with the complexity bounds established for the method DFQRM in grapiglia2024quadratic . However, to take advantage of potential compressibility of gradients, one should take $s_{0}$ sufficiently small such that $m_{0}\ll n$. In this case, the worst-case complexity bound becomes $\mathcal{O}\left(n^{2}\epsilon^{-2}\right)$.